OpenStudy (anonymous):

find the equation of tangent line to the curve x^4+y^4=2 at (1,1)

OpenStudy (anonymous):

y=2-x

OpenStudy (bahrom7893):

4x^3+4y^3(dy/dx)=0 x^3+y^3(dy/dx)=0